Table of Contents

SITE 1167

Hole 1167A (PBF-6A)
The 97-nmi voyage to Site 1167 was accomplished at an average speed of 11.4 kt. The vessel approached the GPS coordinates of the site on 22 February at 2200 hr, and a beacon was deployed at 2238 hr. Hole 1167A was spudded with the APC at 0925 hr on 23 February. The seafloor depth was established from the recovery of the first core at 1651.3 mbsl. APC coring advanced without incident, but with varying recovery (77% to 103%) to 39.7 mbsf. APC refusal was reached when the core barrel of Core 6H did not achieve a full stroke and was recovered empty and partially bent. A successful Adara tool heat-flow measurement was taken at the mudline and on Core 5H. Coring with the XCB system resumed at 1445 hr with Core 7X and continued through Core 21X to a depth of 179.2 mbsf, when operations had to be suspended because of an approaching iceberg. After pumping a 30-bbl sepiolite mud sweep, the pipe was pulled to 42.6 mbsf while the movement of the iceberg was monitored. The iceberg came to within 0.1 nmi before sliding past the drill site. The decision was made to run back to bottom and resume coring by 1000 hr, when the iceberg had reached a range of 0.3 nmi and was moving away from the drill site. XCB coring was resumed and continued until Core 28X was recovered from a depth of 246.5 mbsf. Yet another iceberg approached the drill site at 2345 hr on 24 February. This iceberg came to within 0.5 nmi of the vessel. The iceberg was monitored for an hour, during which time it passed the drill site and was moving away. The drill string was again run to bottom with the top drive still in place. At a depth of 161.7 mbsf the driller noted 25,000 lb of downward drag. Light reaming was required to reach bottom. A core barrel was deployed, and at 0600 hr on 25 February XCB coring resumed.

Coring operations were short lived, however, when another iceberg arrived on the scene. After recovering Core 34X from a depth of 303.2 mbsf, the drill string was once again pulled back to the seafloor, with the end of the pipe placed at 42.6 mbsf.

The plan was to deploy a FFF once the pipe was at a safe depth below the seafloor. This plan had to be changed when the iceberg increased its approach speed and changed course, moving directly toward Site 1167. The decision was made to not deploy the FFF and to remain in position to pull the remaining drill pipe free of the seafloor should the need occur.

On 25 February at 2145 hr, the iceberg had moved to a distance of 1.5 nmi from the drill site and was continuing away at a rapid rate. The decision was made to deploy the FFF at this point because more icebergs were in the vicinity. The drill string was once again run into the hole. As before, the driller encountered an obstruction at 168.7 mbsf, resulting in 25,000 lb of down drag and requiring light reaming to reach the bottom.

Coring continued to Core 49X to a total depth of 447.5 mbsf. Recovery and rate of penetration were extremely variable throughout the coring cycle as we intermittently drilled through coarse sand and gravel beds and encountered occasional dropstones. The decision was made to halt coring operations short of the 620 mbsf objective to conserve adequate time for wireline logging and LWD operations, and to meet the 15-hr early departure from the site required as an additional transit-time contingency.

Logging Operations in Hole 1167A
In preparation for logging, the XCB bit was placed at a depth of 86.9 mbsf, and the Schlumberger wireline sheaves were rigged up. The first suite of logging tools to be deployed was the triple combo consisting of the DITE, HLDS, APS, and HNGS. The tools were deployed on 27 February at 1905 hr; however, they could only be lowered to a depth of 148.7 mbsf, or 61.7 m below the end of the pipe. This short section was logged back to the bit. The tool could not be retracted back into the pipe without circulating the rig pumps to open the flapper valve.

While the logging tools were being recovered, it was decided that further wireline logging efforts had to be abandoned. The logging tools were laid out, and by 2400 hr on 27 February the Schlumberger wireline sheaves were rigged down. The hole was filled with a 30-m cement plug, and the pipe was pulled clear of the seafloor by 0215 hr. The bit cleared the rotary table at 0825 hr, ending Hole 1167A. Total time lost because of icebergs and ice-related problems amounted to 26.75 hr at Hole 1167A.

Hole 1167B
The vessel was offset 50 m to the northwest for a dedicated M/LWD hole. After waiting on weather for 4.25 hr, the drill string was tripped to the bottom and Hole 1167B was spudded at 2100 hr on 28 February. Drilling with the M/LWD system proceeded smoothly throughout the night with excellent results. The time allocated for this operation ran out at 0930 hr on 29 February, and drilling was halted at a depth of 261.8 mbsf. All M/LWD systems and the real time data telemetry equipment performed perfectly.

The hole was displaced with 66 bbl of bentonite gel mud, and the pipe was recovered, clearing the seafloor at 1110 hr on 29 February. During the pipe trip, the positioning beacons were released and recovered. While we attempted to release the third beacon, the portable command unit cable was sucked into the No. 6 thruster well. The cable was severed immediately, and the transducers head was lost. As a result beacon No. 3 could not be released. The hydrophones and thrusters were retracted, and the drilling equipment was secured for transit. At 1730 hr on 29 February 2000, the vessel departed the last site of Leg 188.

Table of Contents